The magnetic induction at a point P which is at a distance of 4cm from a long current - carrying wire is 10 to power -3 T. The field of induction at a distance of 12cm from the current wire will be

Grade:10

1 Answers

Arun
5 years ago
We know B=10^-7*2I/r where B =magnetic field of a long current carrying conductor
I=amount of current 
r=distance from current element
10^-3=10^-7*2*I/4*10^-3
=>I=20 A
again B=10^-7*2*20/12*10^-3=
3.33*10^-4